OpenStudy (anonymous):

3sinx-1=cos 2x for (0 less is or equal to x which is less than or equal to pi) All angles are in radians.

OpenStudy (anonymous):

\[0\le x \le \pi \]

OpenStudy (anonymous):

Using \[\cos(2x) = 1-2\sin^2(x)\]Let y = sin(x), solve the quadratic for y then back for x.

OpenStudy (anonymous):

Okay....soo would my answers be pi/6 and 5pi/6?

OpenStudy (anonymous):

yeah it is okay. :) awesome. Thanks!